### v2

# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections has achieved state-of-the-art energy precision in solving Fermi-Hubbard-type models tensorbackflow . However, the Fermi-Hubbard model is very challenging to solve, and the validity of a method relies on investigating the ground state’s physical property. For simplicity, we name the tensor representation of backflow corrections as the Tensor-Backflow in this work. We apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices as large as 256 sites, under various interaction strengths $U$, electron fillings $n$ and boundary conditions. Energy precision can be further improved by considering more backflow terms, such as considering backflow terms from next-nearest-neighbours or from all sites. Energy extrapolations on 64-site lattices give competitive results to the gradient optimized fPEPS with the bond dimension as large as $D$=20. For cases of $n$=0.875 and $U$=8 on the $16\times 16$ lattice under open boundary condition, by considering nearest-neighbour backflow terms, obtained energy is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with the bond dimension $D$=20. For periodic boundary condition, the variational wave-function is not enforced on any prior symmetry, meanwhile linear stripe order is successfully obtained. Under the same filling and $U$=10,12, energies obtained from initializations with the obtained wave-function for $U$=8 are lower than that from direct optimizations, meanwhile energies are competitive to or even better than state-of-the-art tensor network methods. For cases of $n$=0.8 and 0.9375, results consistent with the phase diagram from AFQMC are obtained by direct optimizations.
I Introduction
The Fermi-Hubbard model is a simple quantum lattice model for describing correlated electrons. Solving the Fermi-Hubbard model is important for understanding the mechanism of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, investigating the Fermi-Hubbard model on a finite-sized lattice is very challenging. Because states with different orders near the ground state have energies so similar that the optimization is difficult. For finite systems, the lattice size should be sufficiently large to reduce the boundary effect. Therefore, to identify the property of the ground state, numerical methods are required to be unbiased and scalable. To achieve this goal, many numerical methods have been developped and benchmarked Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
For example, the density-matrix-renormalization-group (DMRG) is suitable for one-dimensional or quasi one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but the accuracy is not satisfactory for two-dimensional systems. The Quantum-Monte-Carlo (QMC) has very high accuracy, but computational complexity is too high for systems within the “sign problem” QMC_sign . The projected-entangled-pair-state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimizations tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , however the computational complexity is high especially for periodic boundary condition (PBC). For the infinite-PEPS (iPEPS) iPEPS and the density-matrix-embedding-theory (DMET) Hubbard_DMET , the size of the supercell introduces a prior condition on the ground state.
The variational-monte-carlo (VMC) has no “sign problem”, however its effectiveness depends on the variational wave-function. Recently, the neural-network based variational wave-functions have been intensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . By the deep structure, neural-networks can capture long-range correlations or entanglements in the quantum state, thus achieving state representation ability beyond the area-law entanglement NN_represent1 ; NN_represent2 . However, the complex structure of neural-networks introduces difficulties for optimizations. For example, although the backflow corrections by neural-networks have obtained high accuracy in continuous systems NN_continuous1 ; NN_continuous2 , energy precision for the latticed Fermi-Hubbard model is not satisfactory on large lattices Hubbard_NN1 ; Hubbard_NN2 . Recently, the Tensor-Backflow method has achieved state-of-the-art energy precision at reasonable optimization costs for Fermi-Hubbard-type models tensorbackflow . The wave-function is based on backflow corrections backflow1 ; backflow2 , meanwhile the representation ability is significantly enhanced by a way to increase the variational parameter number via embedding independent variables into independent degrees of freedoms into a tensor.
In this work, we use the Tensor-Backflow along with one Lanczos optimization lanczos to investigate the Fermi-Hubbard model under various electron fills $n$, interaction strengths $U$ and boundary conditions on lattices as large as 256 sites. Without prior enforcements on symmetries, energies obtained are competitive to or even better than current state-of-the-art results, meanwhile energy precision can be further improved by considering more backflow terms. For $n$=0.875 and $U$=8 on the $16\times 16$ lattice, ground state with a linear stripe order is obtained by a direct optimization. Periods for spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For the same filling and $U$=10,12, achieved energies from the initialization with the wave-function obtained for $U$=8 are lower than those from direct optimzations. For $n=$0.8 and 0.9375, we compare moderate strengths of $U$ and obtain results consistent with the phase diagram from AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
This paper is organized as follows: Sec.II introduces the Tensor-Backflow method as well as the VMC optimization and a Lanczos optimization. Computational complexity of the method is introduced. Subsection.III.1 presents energy comparisons respect to current state-of-the-art methods such as fPEPS, tensor networks and AFQMC. Subsection.III.2 presents results for $n$=0.875, including energy comparisons, ground state patterns, charge and spin structure factors. Subsection.III.3 presents results for $n$=0.8 and 0.9375, with energy results as well as charge and spin structure factors. Sec.IV concludes this paper.
II the Tensor-Backflow method
II.1 Wave-function definitions
Based on literature tensorbackflow , we proposed an unrestricted form of backflow corrections: the summation for backflow terms are not only on positions but also on spins. The fictitious orbital is constructed as
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}).$ (1)
Eq.(1) has limited number of variational parameters thus is not sufficient for state representation. For the lattice model, independent parameters in Eq.(1) are discrete values, thus we consider each independent parameter in the formula as a degree of freedom in the tensor $g$. The tensor $g$ has the dimension tensorbackflow :
$[M,N,d,Q,d],$ (2)
where $M$ is the particle number, $N$ is the site number, $d$ is the degrees of freedom each site and $Q$ is the number of backflow terms including the particle itself. For example, $Q=5,N$ when considering backflow terms from nearest-neighbours and all sites, respectively. We consider coefficients $c_{ij}$ in Eq.(1) depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ instead of $|\mathbf{S}\rangle$ for simplicity backflow1 ; backflow2 . Thus, first and second $d$ in Eq.(2) are for $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. The number of variational parameters is the multiplication of all degrees of freedoms in tensor $g$.
The wave-function is a Slater-determinant of fictitious particles:
$w(\mathbf{S})=\mathrm{det}M^{B},$ (3)
where the matrix element $M_{ik}^{B}$ is constructed based on the tensor representation of Eq.(1):
$M_{ik}^{B}=\sum_{q}g[i,k,\mathbf{s}(\mathbf{r}_{i}),q,\mathbf{s}(\mathbf{r}_{q})],$ (4)
where $q$ is the summation index including the backflow terms and the $i$-th particle position.
II.2 Optimization method
The wave-function is first optimized by the VMC, then further optimized by a Lanczos step. The energy and the $\alpha$-th parameter’s gradient are evaluated through the Markov-Chain-Monte-Carlo (MCMC) process :
$\begin{split}&E=\langle E_{\mathrm{loc}}\rangle\\ &G^{\alpha}=2\langle E_{\mathrm{loc}}O_{\mathrm{loc}}^{\alpha}\rangle-2\langle E_{\mathrm{loc}}\rangle\langle O_{\mathrm{loc}}^{\alpha}\rangle,\end{split}$ (5)
[…]
The variational parameters are updated according to the gradient descent method. Here we only adopt the first-order gradient descent due to the low optimization difficulty of the tensor representation. Because of the limited MC sample number, we take the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. Such parameter updating scheme has been successfully used in optimizing high dimensional tensors like the PEPS .
A Lanczos step further improves the representation ability of wave-function $|\Psi_{p=0}\rangle$ by considering an additional wave-function $|\Psi_{p=0}^{\perp}\rangle$ orthogonal to $|\Psi_{p=0}\rangle$ lanczos :
$|\Psi_{p=1}\rangle=A|\Psi_{p=0}\rangle+B|\Psi_{p=0}^{\perp}\rangle,$ (6)
where $A$ and $B$ are parameters to be determined, and $|\Psi_{p=0}\rangle$ is the wave-function obtained after the VMC. The orthogonal wave-function is built by $|\Psi_{p}^{\perp}\rangle=(\hat{H}-E_{p})/\sigma_{p}|\Psi_{p}\rangle$, where the energy expectation $E_{p}=\langle\Psi_{p}|\hat{H}|\Psi_{p}\rangle$ and the variance $\sigma_{p}^{2}=\langle\Psi_{p}|(\hat{H}-E_{p})^{2}|\Psi_{p}\rangle$.
II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity for generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$ based on Eq.(3) and Eq.(4), where $\mathcal{O}(N^{3})$ is from the Slater determinant and $\mathcal{O}(Q)$ is from the summation in Eq.(4). When considering backflow terms from nearest-neighbour sites, there is $Q$=5, meanwhile the complexity is $\mathcal{O}(N^{3})$. The maximal representation ability is achieved by considering backflow terms from all sites: $Q=N$, meanwhile the complexity is $\mathcal{O}(N^{4})$. For the wave-function after a Lanczos optimization, based on Eq.(6), the complexity is $\mathcal{O}(N^{4}Q)$ for generating a coefficient $w(\mathbf{S})$.
The interval between two MC samples is $N$ during the optimization, thus the total complexity for one VMC step is $\mathcal{O}(N^{4}Q)$. In VMC optimizations, sample number and step size are important for achieving fully convergences. To identify these parameters, we have tested two optimization strategies: (1)under a particular sample number in the magnitude of $\mathcal{O}(10^{4})$, firstly setting the step size as $\mathcal{O}(10^{-3})$ and then reducing it to $\mathcal{O}(10^{-4})$; (2)under the sample number $\mathcal{O}(10^{4})$, using the step size $\mathcal{O}(10^{-3})$, then increasing the sample number to $\mathcal{O}(10^{5})$, lastly reducing the step size to $\mathcal{O}(10^{-4})$. We found that converged energies achieved by both strategies only have differences on the fourth decimal place. For both strategies, $\mathcal{O}(1\times 10^{4})$ optimization steps are applied for fully convergence. In the Lanczos optimization, the complexity is $\mathcal{O}(N^{5}Q)$ due to the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ when calculating the variance $\sigma_{p}^{2}$.
III Numerical results
In this section, we present numerical results for the two-dimensional Fermi-Hubbard model achieved by the Tensor-Backflow method. To demonstrate the method’s high accuracy, we firstly present energy comparisons respect to previous state-of-the-art results.
The Hamiltinian of the Fermi-Hubbard model is:
$\begin{split}\hat{H}=&-t\sum_{\langle ij\rangle,\sigma}(\hat{c}_{i\sigma}^{\dagger}\hat{c}_{j\sigma}+\mathrm{H.c.})-t^{\prime}\sum_{\langle\langle ij\rangle\rangle,\sigma}(\hat{c}_{i\sigma}^{\dagger}\hat{c}_{j\sigma}+\mathrm{H.c.})\\ &+U\sum_{i}\hat{n}_{i\uparrow}\hat{n}_{i\downarrow},\end{split}$ (7)
where $t$ and $t^{\prime}$ are hopping strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set $t$=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos optimization are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.
Figure 1: Energy extrapolations and comparisons between Tensor-Backflow and fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). Energy extrapolation based on Tensor-Backflow results under PBC for the lattice of $4\times 16$(c). $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos optimization.
III.1 Comparisons with previous results
Fig.(1)(a)(b) depict energy comparisons respect to the fPEPS method under open-boundary condition (OBC) on 64-site lattices. The fPEPS is a state-of-the-art method for latticed fermions and its state representation ability can be improved by increasing the bond-dimension ($D$). From literature Hubbard_fPEPS , the complexity of fPEPS scales rapidly respect to $D$: $\mathcal{O}(D^{6})$, and the maximal bond dimension is $D$=20. The fPEPS is firstly optimized by the imaginary-time-evolution method named as the simple-update (SU), followed by the gradient-optimization (GO). The fPEPS with $D$=20 can achieve extremely high energy precision under OBC.
In our optimizations, the initial state of Tensor-Backflow is unrestricted Hartree-Fock (UHF) state represented by a tensor of the dimension $[M,N,2]$, where $M$ is the site number $N$ is the particle number and 2 is because of spin. From the initial state, we expand the tensor’s dimension to Eq.(2) to include backflow terms tensorbackflow . To avoid any bias introduced by UHF, optimizations of UHFs are ceased when the energy is far from fully convergence UHF .
In Fig.(1)(a)(b), red and blue solid dots denote NN and all-site backflow terms, respectively. High and low energy values for same-colored solid dots denote without and with one Lanczos optimization, respectively. Therefore, a Lanczos optimization significantly improves energy precision for either NN backflow and all-site backflow. Comparing obtained energies after one Lanczos optimization and the fPEPS with $D$=20, the NN backflow can achieve the energy precision competitive to the SU optimized fPEPS, meanwhile the all-site backflow can achieve the energy precision between SU and GO optimized fPEPS. The Tensor-Backflow has no limitations on boundary conditions. Fig.(1)(c) depicts energy results on the $4\times 16$ lattice under PBC.
To obtain the energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . In Fig.(1), energy extrapolations are denoted by blue solid lines. From Fig.(1)(a)(b), extrapolated energies under OBC for $8\times 8$ is -0.6993, and for $4\times 16$ is -0.6826. Extrapolated energies are close to state-of-the-art energies: -0.6993 for $8\times 8$ and -0.6830 for $4\times 16$, achieved by the fPEPS with $D$=20 within GO optimized. Based on results of OBC, the extrapolated energy of PBC in Fig.(1)(c) predicts the state-of-the-art result: -0.7642.
  $n$  $U$
Lattice
Size
Boundary
Condition
  $E_{p=0}$    $E_{p=1}$    Ref.
0.875 8 $8\times 8$ OBC -0.6955 -0.6982 -0.6993 Hubbard_fPEPS
$4\times 16$0.6797 -0.6821 -0.6830 Hubbard_fPEPS
$16\times 16$0.7219 -0.7257 -0.7260 Hubbard_fPEPS
$16\times 16$ PBC -0.7509 -0.7552 -0.7544 Manybody_benchmark
10 $16\times 16$ A–PBC -0.6786 -0.6828 -0.6813 Hubbard_TNS
1 8 $12\times 12$ PBC -0.5209 -0.5224 -0.5246 AFQMC_n1
A–PBC -0.5205 -0.5224 -0.5249 AFQMC_n1
44/64 4 $8\times 8$ PBC -1.1833 -1.1846 -1.1867 AFQMC_sym
80/100 $10\times 10$1.1086 -1.1109 -1.1147 AFQMC_sym
116/144 $12\times 12$1.1038 -1.1061 -1.1102 AFQMC_sym
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos optimization, respectively.
Energy comparisons are detailed in Tab.(1). In the table, $E_{p=0}$ is from the Tensor-Backflow and $E_{p=1}$ is from a Lanczos optimization of the converged wave-function. For cases under OBC, reference energies are from GO optimized fPEPS with the largest bond dimension $D$=20 Hubbard_fPEPS . For lattices of $8\times 8$ and $4\times 16$, $E_{p=0}$ is achieved by the Tensor-Backflow with all-site backflow terms. For the lattice of $16\times 16$, $E_{p=0}$ is achieved by the Tensor-Backflow with NN backflow terms. $E_{p=1}$ are competitive to fPEPS results, and relative errors repsect to fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$ and $4.1\times 10^{-4}$ for $8\times 8$, $4\times 16$ and $16\times 16$ lattices, respectively. Therefore, comparing to the fPEPS, energy precision for NN backflow on the $16\times 16$ lattice is even better than that for all-site backflow on 64-site lattices.
Because of the high complexity, cases under PBC are not practical for fPEPS. Reference energy for $n$=0.875 and $U$=8 on the $16\times 16$ lattice under PBC is -0.7544 from the VMC method in literature Manybody_benchmark . Meanwhile, the Tensor-Backflow after a Lanczos optimization has achieved the energy $1\times 10^{-3}$ lower. Under the same filling and $U$=10, with antiperiodic along the x-boundary and periodic along the y-boundary (A-PBC in the table), the Tensor-Backflow after a Lanczos optimization has achieved the energy $2.2\times 10^{-3}$ lower comparing to the reference energy. The reference energy is from the energy extrapolation result of the tensor network method in literature Hubbard_TNS . Comparing to AFQMC results in literatures AFQMC_n1 ; AFQMC_sym , the Tensor-Backflow after a Lanczos optimization has achieved competitive energies under various fillings, interaction strengths and boundary conditions. Meanwhile, relative errors are all in the magnitude of $\mathcal{O}(10^{-3})$.
Backflow
Type
Lattice
Size
Parameter
Number
Relative
Error(p=0)
Relative
Error(p=1)
NN $8\times 8$ 286720 $6.7\times 10^{-3}$ $2.3\times 10^{-3}$
$4\times 16$ 286720 $8.0\times 10^{-3}$ $3.6\times 10^{-3}$
all-site $8\times 8$ 3670016 $5.4\times 10^{-3}$ $1.6\times 10^{-3}$
$4\times 16$ 3670016 $4.9\times 10^{-3}$ $1.3\times 10^{-3}$
NN $16\times 16$ 4587520 $5.6\times 10^{-3}$ $4.5\times 10^{-4}$
Table 2: Comparisons between parameter number used in Tensor-Backflow for various site number and backflow types, with $n$=0.875 and $U$=8 under OBC. Reference energies are from the GO optimized fPEPS with the bond dimension $D$=20. Parameter number is based on Eq.(2), where $Q=5,N$ for NN and all-site backflow, respectively.
Tab.(2) compares parameter number used in Tensor-Backflow for various lattice sizes and backflow types. According to Eq.(2), for the filling of $n$=0.875 on 64-site lattices, the parameter number for NN backflow and all-site backflow are 286720 and 3670016, respectively. Meanwhile, the parameter number for NN backflow on the $16\times 16$ lattice is 4587520. Although the parameter number for NN backflow on the $16\times 16$ lattice and all-site backflow on 64-site lattices are in the same magnitude, the relative error on the $16\times 16$ lattice is much lower. Because only NN hopping terms in the Fermi-Hubbard model are considered, thus the primary state representation is achieved by NN backflow terms, meanwhile the improvement of all-site backflow is moderate. For NN backflows, the parameter number grows as $\mathcal{O}(N^{2})$. When representing the Hubbard model with NN hoppings, a large parameter number of $4.6\times 10^{6}$ on the $16\times 16$ lattice has better state representation ability comparing to the small parameter number of $2.9\times 10^{5}$ on 64-site lattices.
On the $16\times 16$ lattice under OBC, the fPEPS with $D$=14 has $5.7\times 10^{6}$ parameters, and the achieved energy is -0.7224. The fPEPS with $D$=16 has $9.4\times 10^{6}$ parameters, and the achieved energy is -0.7251. The fPEPS with $D$=18 has $1.5\times 10^{7}$ parameters, and the achieved energy is -0.7260. The fPEPS with the maximal bond dimension $D$=20 has $2.2\times 10^{7}$ parameters, with the similar energy as $D$=18 Hubbard_fPEPS . For the Tensor-Backflow on the $16\times 16$ lattice, there are $E_{p=0}=$-0.7219 and $E_{p=1}=$-0.7257. Under the similar energy precision, without a Lanczos optimization, the Tensor-Backflow has many lower parameters comparing to the fPEPS with $D$=14. After a Lanczos optimization, same energy precision is achieved by the Tensor-Backflow with significantly lower parameters comparing to the fPEPS with $D$=18,20. Therefore, the Tensor-Backflow is an efficient state representation.
III.2 n=0.875
The filling of $n$=0.875 is important for understanding the mechanism of superconductivity, meanwhile the ground energies are highly competitive between various methods Hubbard_stripe .
Figure 2: Patterns of ground states for $n$=0.875, $U$=8 under PBC(a)(b) and OBC(c)(d). Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$.
On the $16\times 16$ lattice, a direct optimization for $U$=8 under PBC achieves the linear stripe order, and patterns of the ground state are depicted in Fig.(2)(a)(b). Fig.(2)(a) depicts the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$ and Fig.(2)(b) depicts the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. As depicted by the ground state’s pattern, periods for the SDW and CDW are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational invariance is not enforced in the variational wave-function, even the obtained energy is competitive, the optimization is converged to a subplace of the true ground state.
For the same filling $n$ and the interaction strength $U$, we used the obtained wave-function under PBC as the initial state and then performed optimizations under OBC. The achieved state-of-the-art energy is -0.7257 as reported in Tab.(1) and (2). Meanwhile, Fig.(2)(c)(d) depict the average spin and charge density per site, respectively. Even the stripe order in the initial state has the same period as the ground state under OBC, the optimization can slightly modify the stripe’s pattern according to the boundary condition. Patterns of the ground state under OBC yield to fPEPS results Hubbard_fPEPS .
Lattice
Size
Boundary
Condition
  $U$
Direct
Optimization
Initialization
with $U$=8
$16\times 16$ PBC 8 -0.7552
10 -0.6815 -0.6831
12 -0.6303 -0.6317
A-PBC 10 -0.6813 -0.6828
Table 3: Energy comparisons between different initializations for the filling $n$=0.875. The direct optimization of $U$=8 achieves the ground state of linear stripe order, and such order is preserved when it is the initial state for other $U$ and boundary conditions. Energies with linear stripe orders are much lower than those by direct optimizations.
Figure 3: Charge and spin structure factors for the filling $n$=0.875 under various $U$ and boundary conditions. (a)(b)On the $16\times 16$ lattice, results of $U$=8 under PBC are obtained by a direct optimization, meanwhile other results are obtained by the initialization with the $U$=8 PBC wave-function. (c)(d)Results of $U$=8 under PBC by a direct optimization on the $8\times 32$ lattice.
For other interaction strengths such as $U$=10,12, Tab.(3) denote energy comparisons between the linear stripe order and direct optimizations, where linear stripe order is preserved when initializing the optimization with the wave-function obtained for $U$=8 and PBC. All energies are evaluated by $p$=1 wave-functions. In particular, direct optimization achieves diagonal stripe order when $U$=12 under PBC, and direct optimizations achieve linear SDW however no obvious CDW when $U$=10 under both PBC and A-PBC. However, the energy of linear stripe order is much lower than that of the direct optimization. Therefore, it’s concluded that the ground state is linear stripe order for large interaction strengths of $U$.
Lattice
Size
   $t^{\prime}$
Backflow
Type
   $E_{p=0}$    $E_{p=1}$
$8\times 16$ 0 NN -0.7529 -0.7568
$8\times 24$0.7500 -0.7546
$8\times 32$0.7505 -0.7547
$8\times 16$0.2 -0.7354 -0.7406
$8\times 16$ NNN -0.7359 -0.7410
$12\times 12$0.7333 -0.7388
$16\times 16$0.7335 -0.7386
Table 4: Energy results obtained by direct optimizations for the filling $n$=0.875 and $U$=8. $t^{\prime}$ is defined in Eq.(7). For non-zero $t^{\prime}$, NNN backflow terms are considered in the Tensor-Backflow. On the $8\times 16$ lattice, considering NNN backflow terms achieves the energy slightly lower than that when only considering NN backflow terms.
Figure 4: Patterns of ground states for $n$=0.875, $U$=8 and $t^{\prime}$=-0.2 under PBC on the $8\times 16$ lattice. Left-sided (right-sided) figure denotes the average spin (density) per site, within definitions from Fig.(2).
Figure 5: Comparisons of spin structure factors between $t^{\prime}$=0 and -0.2 for $U$=8 under PBC. Number of k-points are increased from $17^{2}$ to $65^{2}$ when calculating spin structure factors. The horizontal axis is $k_{x}$ for $t^{\prime}$=-0.2 and $k_{y}$ for $t^{\prime}$=0, meanwhile vertical axis is $S(k_{x},k_{y}=\pi)$ for $t^{\prime}$=-0.2 and $S(k_{x}=\pi,k_{y})$ for $t^{\prime}$=0.
Charge and spin structure factors determine periods of the stripe order in the ground state. The structure factors are defined as: $C(\textbf{k})=\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(2). $C(\textbf{k})$ is set to be zero when $k_{x}=k_{y}$=0. All results of structure factors are obtained by $p$=1 wave-functions. Fig.(3) depicts structure factors for the filling of $n$=0.875 under various $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. A direct optimization for $U$=8 under PBC achieves the linear stripe order. When applies the wave-function as the initial state for $U$=10,12 under various boundary conditions, linear stripe orders are preserved. As depicted by Fig.(3)(a)(b), charge structure factor peaks at $\pi/4$ and spin structure factor peaks at $87\pi/$ 8, corresponding to periods of 8 and 16 for CDW and SDW, respectively. Fig.(3)(c)(d) depicts results of a direct optimization on the $8\times 32$ lattice for $n$=0.875 and $U$=8 under PBC. A vertical stripe order is obtained, meanwhile structure factors peaks at same locations as on the $16\times 16$ lattice. On the $16\times 16$ lattice, the lattice’s side length of 16 corresponds to a whole period of the SDW, thus peaks on the $16\times 16$ lattice are higher than those on the $8\times 32$ lattice.
Tab.(4) depict additional energy results for $n$=0.875 and $U$=8 under PBC. All results in the table are obtained by direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, the vertical stripe order is obtained with the periods same as those on the $16\times 16$ lattice. Specifically, we investigate the case for non-zero $t^{\prime}$, which is promising to have superconductivity Hubbard_SC1 . For $t^{\prime}=-0.2$, we find that NNN backflow achieves slightly lower energy than that by NN backlfow, on the $8\times 16$ lattice. Thus, NNN backflow terms are considered for non-zero $t^{\prime}$.
  $n$   $U$
Lattice
Size
Backflow
Type
  $E_{p=0}$    $E_{p=1}$
0.8 4 $10\times 20$ NN -1.1012 -1.1037
8 $10\times 20$0.8755 -0.8793
$10\times 10$0.8801 -0.8848
$10\times 10$ all-site -0.8809 -0.8853
0.9375 4 $8\times 16$ NN -0.9384 -0.9409
$8\times 24$0.9394 -0.9421
$8\times 32$0.9401 -0.9426
Table 5: Energy results obtained by direct optimizations for the filling $n$=0.8 and $n$=0.9375, with $t^{\prime}$=0 under PBC. On the $10\times 10$ lattice, all-site backflow has moderate improvement comparing to NN backflow for $n$=0.8 and $U$=8.
Figure 6: Charge and spin structure factors on the $10\times 20$ lattice for $n$=0.8 (a)(b) and on the $8\times 32$ lattice for $n$=0.9375 (c)(d), under PBC. (a)(b)For $n$=0.8, the spin structure factor peaks at $54\pi/$ 5 when $U$=8, however there are no obvious peaks when $U$=4. For both $U$=4,8, there are no obvious peaks in the charge structure factor. (c)(d)For $n$=0.9375, the spin structure factor peaks at $1615\pi/$ 16 when $U$=4, meanwhile there are no obvious peaks in the charge structure factor.
Fig.(5) depicts the ground state pattern for $t^{\prime}$=-0.2 on the $8\times 16$ lattice, with definitions of average spin and density per site from Fig.(2). From the figure, an obvious width-4 stripe is achieved, which yields to the recent AFQMC results on width-8 lattices Hubbard_SC1 . From results on larger lattices such as $12\times 12$ and $16\times 16$, there are no obvious patterns from $s_{i}^{z}$ or $n_{i}$. In Fig.(5), we compare spin structure factors between $t^{\prime}$=0 and -0.2. To reval detailed differences, we increase the number of k-points from $17^{2}$ to $65^{2}$. Thus, there are totally 33 k-points in the x-axis. From the figure, for $t^{\prime}$=0 peaks are sharply distrubuted and centered on $87\pi/$ 8. However, the peak for $t^{\prime}$=-0.2 is wider comparing to that for $t^{\prime}$=0. The wider peak reveals competitions between stripe orders of different periods in the ground state for $t^{\prime}$=-0.2, thus direct optimizations are easier to obtain local optimizations on larger lattices.
III.3 n=0.8 and n=0.9375
Energy results for $n$=0.8 and $n$=0.9385 are presented on Tab.(5). As depicted by cases of $n$=0.8, $U$=8 on the $10\times 10$ lattice, the improvement of all-site backflow is moderate comparing to NN backflow. Thus, the majority improvement of the state representation is achieved by NN backflow terms.
Fig.(6) depicts charge and spin structure factors for $n$=0.8 on the $10\times 20$ lattice and for $n$=0.9375 on the $8\times 32$ lattice. Based on the phase diagram via AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ between un-ordered and ordered ground states increase with the doping level $h$. Here, for $n$=0.8, the spin structure factor peaks at $54\pi/$ 5 for $U$=8, meanwhile there are no obvious peaks for $U$=4. Comparing charge structure factors between $U$=4 and 8, further increasing $U$ will lead to the obvious peak in the charge structure factor. For $n$=0.9375, the spin structure factor peaks at $1615\pi/$ 16. Results of $n$=0.8 and 0.9375 obtained by direct optimizations yield to the phase diagram via AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
IV Discussions and conclusions
Figure 7: Density-density correlations for $n$=0.875, $U$=8 on the $8\times 32$ lattice under PBC. Blue dots depict the $p$=0 wave-function and yellow dots depict the $p$=1 wave-function. For both $p$=0 and $p$=1, results are obtained based on 46080 MC samples, with the interval between two samples twice as the lattice size.
In this work, we benchmarked the Tensor-Backflow method along with a Lanczos optimization for various fillings $n$, interaction strengths $U$, NNN strengths $t^{\prime}$, lattice sizes and boundary conditions. We investigate ground state properties through charge and spin structure factors, and results consistent with other state-of-the-art methods such as fPEPS for $n$=0.875 as well as AFQMC for $n$=0.8 and 0.9375 are obtained.
From our results, although a Lanczos optimization significantly improves the energy, the property of the state is determined by the $p$=0 wave-function. Fig.(7) depicts density-density correlations from $p$=0 and $p$=1 wave-functions for $n$=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. From the figure, the CDW with the period of 8 is already in the $p$=0 wave-function and there are no significant differences for $p$=1 results. Thus, we evaluate pair-pair correlations with definitions from literature pair_pair_correlation based on $p$=0 wave-functions for the case of $n$=0.875 and $U$=8 on the $16\times 16$ lattice. For either $t^{\prime}$=0 and -0.2, pair-pair correlations $P(d)$ are in the magnitude of $\mathcal{O}(10^{-5})$ when $d_{x}$=8 or $d_{y}$=8. Because energies with different orders are competitive when $n$=0.875, obtaining local optimizations are common under direct optimizations. Therefore, more careful and intensive investigations are necessary for $n$=0.875.
Based on results in this work, the Tensor-Backflow is an efficient state representation for the latticed Fermi-Hubbard model. Furthermore, there are several ways to improve the energy precision. For example, considering more backflow terms rather than NN backflow to improve energy precision for the Hubbard model with NN hoppings, comparing energies under various initializations, applying more advanced optimization methods such as the stochastic reconfiguration, etc. Currently, the wave-function is not symmetry enforced, thus translational invariance is not preserved during optimizations. For PBC, the energy precision can be further improved by enforcing translational invariance on the wave-function. The Tensor-Backflow has no restrictions on lattice shapes and boundary conditions, thus it is potentially an efficient and universal method for solving latticed fermion systems.
V Acknowledgement
X. Liang thanks useful discussions with S. Zhang, W.-Y. Liu and A. Chen. X. Liang thanks Eric J. Walter for resource allocations on the William & Mary Sciclone cluster. The author acknowledges William & Mary Research Computing and the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the results reported within this paper. Part of numerical results were obtained on the Bridges-2 at Pittsburgh Supercomputing Center through allocation (phy240244).