### v3

# Investigating the Fermi-Hubbard model by the Tensor-Backflow method

Abstract
Recently, a variational wave-function based on the tensor representation of backflow corrections has achieved state-of-the-art energy precision in solving Fermi-Hubbard-type models tensorbackflow . However, the Fermi-Hubbard model is very challenging to solve, and the validity of a method relies on investigating the ground state’s physical property. For simplicity, we name the tensor representation of backflow corrections as the Tensor-Backflow in this work. We apply the Tensor-Backflow method to investigate the Fermi-Hubbard model on two-dimensional lattices as large as 256 sites, under various interaction strengths $U$, electron fillings $n$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions. Instead of considering backflow terms from all sites, competitive results are achieved by considering nearest-neighbor or next-nearest-neighbor backflow terms. Meanwhile the variational wave-function is not enforced on geometric symmetries. When $t^{\prime}$=0, by considering nearest-neighbor backflow terms, linear stripe order is sucessfully obtained for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under periodoc boundary condition. For a similar case under open boundary condition, obtained energy is only $4.5\times 10^{-4}$ higher than the state-of-the-art method fPEPS with the bond dimension DD=20. Comparing to state-of-the-art neural network results, energies are competitive and relative errors are below $5\times 10^{-3}$. For cases of nn=0.8 and 0.9375, results consistent with the phase diagram from AFQMC are obtained by direct optimizations. When $t^{\prime}$=-0.2, by considering next-nearest-neighbor backflow terms, obtained energies are competitive or even lower than state-of-the-art neural network results. For example, obtained energy for nn=0.875, UU=8 on the $12\times 12$ lattice under PBC is $8.1\times 10^{-4}$ lower comparing to that from the neural network state. Therefore, the Tensor-Backflow method has strong representation abilities for the Fermi-Hubbard model.
I Introduction
The Fermi-Hubbard model is a simple quantum lattice model for describing correlated electrons. Solving the Fermi-Hubbard model is important for understanding the mechanism of superconductivity Hubbard_SC1 ; Hubbard_SC2 ; Hubbard_SC3 . However, investigating the Fermi-Hubbard model on a finite-sized lattice is very challenging. Because states with different orders near the ground state have energies so similar that the optimization is easily to get trapped on a local minima. For finite systems, the lattice size should be sufficiently large to reduce the boundary effect. Therefore, to identify the property of the ground state, numerical methods are required to be unbiased and scalable. To achieve this goal, many numerical methods have been developed and benchmarked Hubbard_benchmark ; Manybody_benchmark ; Hubbard_DMET ; Hubbard_DMRG ; Hubbard_NN1 ; Hubbard_NN2 ; Hubbard_NN3 ; Hubbard_NN4 ; Hubbard_TNS ; tJ_fPEPS ; Hubbard_fPEPS .
For example, the density-matrix-renormalization-group (DMRG) is suitable for one-dimensional or quasi one-dimensional systems Hubbard_benchmark ; Hubbard_DMRG ; Hubbard_DMRG2 ; Hubbard_DMRG3 ; Hubbard_DMRG4 , but the accuracy is not satisfactory for two-dimensional systems. The Quantum-Monte-Carlo (QMC) has very high accuracy, but computational complexity is too high for systems within the “sign problem” QMC_sign_1 ; QMC_sign_2 ; QMC_sign_3 . The projected-entangled-pair-state (PEPS) can achieve high accuracy for two-dimensional lattices after variational optimizations tJ_fPEPS ; Hubbard_fPEPS ; PEPS_liu , however the computational complexity is high especially for periodic boundary condition (PBC). For the infinite-PEPS (iPEPS) iPEPS and the density-matrix-embedding-theory (DMET) Hubbard_DMET , the size of the supercell introduces a prior condition on the ground state.
The variational-monte-carlo (VMC) has no “sign problem”, however its effectiveness depends on the variational wave-function. Recently, the neural-network based variational wave-functions have been intensively benchmarked RBM_science ; CNN_J1J2_1 ; CNN_J1J2_2 ; CNN_J1J2_3 ; J1J2_SOTA1 ; J1J2_SOTA2 ; J1J2_SOTA3 ; Hubbard_NN1 ; Hubbard_NN2 . By the deep structure, neural-networks can capture long-range correlations or entanglements in the quantum state, thus achieving state representation ability beyond the area-law entanglement NN_represent1 ; NN_represent2 . Although backflow corrections by neural-networks have obtained high accuracy in continuous systems NN_continuous1 ; NN_continuous2 ; NN_continuous3 For the lattice Fermi-Hubbard model, high optimization difficulties of neural networks lead to unsatisfied energy precision on large lattices Hubbard_NN1 ; Hubbard_NN2 or requirements of pre-conditions on wave-functions Hubbard_NN3 ; Hubbard_NN4 . Recently, the Tensor-Backflow method has achieved state-of-the-art energy precision at reasonable optimization costs for Fermi-Hubbard-type models tensorbackflow . The wave-function is based on backflow corrections backflow1 ; backflow2 , meanwhile the representation ability is significantly enhanced by the tensor representation of the wave-function.
In this work, we use the Tensor-Backflow method well-developed in the previous literature tensorbackflow to investigate the Fermi-Hubbard model under various electron fillings $n$, interaction strengths $U$, next-nearest-neighbor hopping strengths $t^{\prime}$ and boundary conditions on lattices as large as 256 sites. Without prior enforcements on geometric symmetries, achieved energies are competitive or even better than current state-of-the-art results. When $t^{\prime}$=0, the ground state with a linear stripe order is achieved without applying pinning-fields for the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC. Periods for spin-density-wave (SDW) and charge-density-wave (CDW) are 16 and 8, respectively. For cases of nn=0.875, UU=10 and 12, linear stripe orders are preserved when transferring from the wave-function converged for UU=8, meanwhile energies are lower comparing to other orders achieved by direct optimzations. For cases of n=n=0.8 with UU=4, 8 and nn=0.9375 with UU=4, results are consistent with the phase diagram of AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 . For more challenging cases of $t^{\prime}$=-0.2 where non-zero superconductivity is expected Hubbard_SC1 , energy achieved by Tensor-Backflow is competitive or even lower than those from the state-of-the-art neural network state (NQS) Hubbard_NN3 , demonstrating the strong representation ability of Tensor-Backflow.
This paper is organized as follows: Sec.II introduces the Tensor-Backflow method as well as the VMC optimization and a Lanczos optimization. Computational complexity of the method is introduced. Subsec.III.1 compares two optimization strategies. Subsec.III.2 presents energy comparisons respect to current state-of-the-art methods such as fPEPS, AFQMC, NQS, etc. Subsec.III.3 presents results for nn=0.875 with $t^{\prime}$=0 and $t^{\prime}$=-0.2, including energy comparisons, ground state patterns, charge and spin structure factors. Subsec.III.4 presents results for nn=0.8 and 0.9375, with energy results as well as charge and spin structure factors. Sec.IV concludes this paper.
II the Tensor-Backflow method
II.1 Wave-function definitions
In this subsection, we briefly introduce the Tensor-Backflow method developed in the previous literature tensorbackflow . An unrestricted form of backflow corrections is proposed in literature tensorbackflow : the summation for backflow terms are not only on positions but also on spins. The fictitious orbital is constructed as
$\phi^{B}_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})=\phi_{k\sigma_{k}}(\mathbf{r}_{i,\sigma_{i}})+\sum_{j}c_{ij}[\mathbf{S}]\sum_{\sigma_{j}=\pm 1}\phi_{k\sigma_{k}}(\mathbf{r}_{j,\sigma_{j}}),$ (1)
where $c_{ij}[\mathbf{S}]$ is the coefficient depending on the configuration $|\mathbf{S}\rangle$ and $\phi_{k\sigma_{k}}$ is the Hatree-Fock (HF) orbital.
Based on the backflow formula, the NQS can improve the representation ability by learning backflow terms Hubbard_NN1 ; Hubbard_NN3 ; NN_continuous1 ; NN_continuous2 ; NN_continuous3 . Comparing to NQS Hubbard_NN2 ; Hubbard_NN3 , we use tensor representation of the wave-function to improve the representation ability, meanwhile keeping the optimization feasible for first-order-gradient methods tensorbackflow .
For a function with discrete variables of $x_{1}$ and $x_{2}$: $y=f(x_{1},x_{2})$, it can be represented by a tenesor $\xi$: $y=\xi_{x_{1},x_{2}}$, where $x_{1}$ and $x_{2}$ are indices in the $\xi$. We consider coefficients $c_{ij}$ in Eq.(1) depend on local configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$ instead of $|\mathbf{S}\rangle$ for simplicity backflow1 ; backflow2 . Then, the independent variables in Eq.(1) are particle position $\textbf{r}_{i}$, orbital number $k$ for $\sigma_{k}=\pm 1$, backflow indices $j$, configurations $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Based on independent degrees of freedom in the wave-function, a tensor $g$ is constructed with each dimension in $g$ corresponding to the independent variable in Eq.(1) respectively. Therefore, the total dimension of $g$ is tensorbackflow :
$[M,N,d,Q,d],$ (2)
where $M$ is the site number, $N$ is the total particle number, $d$ is the degrees of freedom each site and $Q$ is the number of backflow terms including the particle itself. First and second $d$ are for $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$, respectively. When considering backflow terms from nearest-neighbors and all sites, there is: $Q=5$ and $N$, respectively.
The wave-function is a Slater-determinant of fictitious particles:
$w(\mathbf{S})=\mathrm{det}M^{B},$ (3)
where the matrix element $M_{ik}^{B}$ is constructed based on the tensor representation of Eq.(1):
$M_{ik}^{B}=\sum_{q}g[\mathbf{r}_{i},k,\mathbf{s}(\mathbf{r}_{i}),q,\mathbf{s}(\mathbf{r}_{q})],$ (4)
where $q$ is the summation index including the backflow terms and the $i$-th particle position.
Therefore, the wave-function $w(\mathbf{S})$ is generated by a summation of elements from tensor $g$ based on the backflow rule. In the wave-function, there are no enforcements on translational or rotational symmetries, and the total number of variational parameters is the multiplication of all indices of $g$. Comparing to representing correlations by decomposed tensors CP_tensor , in this work, the representation ability is achieved by the tensor $g$ natually containing high order correlations tensorbackflow . For example, $g$ can represent high order correlations between $\mathbf{s}(\mathbf{r}_{i})$ and $\mathbf{s}(\mathbf{r}_{j})$. Such flexible representation ability of $g$ as well as the simple construction of the wave-function lead to low optimization difficulty, therefore competitive precision is achieved within the first-order-gradient method.
II.2 Optimization method
The wave-function is firstly optimized by the VMC, then further optimized by a Lanczos step. The energy and the $\alpha$-th parameter’s gradient are evaluated through the Markov-Chain-Monte-Carlo (MCMC) process Hubbard_fPEPS ; PEPS_liu :
$\begin{split}&E=\langle E_{\mathrm{loc}}\rangle\\ &G^{\alpha}=2\langle E_{\mathrm{loc}}O_{\mathrm{loc}}^{\alpha}\rangle-2\langle E_{\mathrm{loc}}\rangle\langle O_{\mathrm{loc}}^{\alpha}\rangle,\end{split}$ (5)
where the local energy is $E_{\mathrm{loc}}(\mathbf{S})=\sum_{\mathbf{S}^{\prime}}\frac{w(\mathbf{S}^{\prime})}{w(\mathbf{S})}\langle\mathbf{S}^{\prime}|\hat{H}|\mathbf{S}\rangle$, the $O_{\mathrm{loc}}(\mathbf{S})^{\alpha}=\frac{1}{w(\mathbf{S})}\frac{\partial w(\mathbf{S})}{\partial\alpha}$, and $\langle\cdots\rangle$ denotes the average on MCMC samples.
The variational parameters are updated according to the gradient descent method. Here we only adopt the first-order gradient descent due to the low optimization difficulty of the tensor representation. Because of the limited MC sample number, we take the sign of the gradient and apply a constant step size $\delta$: $\alpha^{\prime}=\alpha-\delta\mathrm{sgn}(G^{\alpha})$. Such parameter updating scheme has been successfully used in optimizing high dimensional tensors like the PEPS tJ_fPEPS ; PEPS_liu .
A Lanczos step further improves the representation ability of wave-function $|\Psi_{p=0}\rangle$ by considering an additional wave-function $|\Psi_{p=0}^{\perp}\rangle$ orthogonal to $|\Psi_{p=0}\rangle$ lanczos :
$|\Psi_{p=1}\rangle=A|\Psi_{p=0}\rangle+B|\Psi_{p=0}^{\perp}\rangle,$ (6)
where $A$ and $B$ are parameters to be determined, and $|\Psi_{p=0}\rangle$ is the wave-function obtained after the VMC. The orthogonal wave-function is built by $|\Psi_{p}^{\perp}\rangle=(\hat{H}-E_{p})/\sigma_{p}|\Psi_{p}\rangle$, where the energy expectation $E_{p}=\langle\Psi_{p}|\hat{H}|\Psi_{p}\rangle$ and the variance $\sigma_{p}^{2}=\langle\Psi_{p}|(\hat{H}-E_{p})^{2}|\Psi_{p}\rangle$.
II.3 Computational complexity
For the Tensor-Backflow wave-function, the complexity for generating a coefficient $w(\mathbf{S})$ is $\mathcal{O}(N^{3}Q)$ based on Eq.(3) and Eq.(4), where $\mathcal{O}(N^{3})$ is from the Slater determinant and $\mathcal{O}(Q)$ is from the summation in Eq.(4).
When considering backflow terms from nearest-neighbor (NN) sites, there is QQ=5. The representation ability can be improved by considering backflow terms in the range of next-nearest-neighbor (NNN) sites, there is QQ=9. We denote the above two types as NN and NNN backflows, and the complexities are $\mathcal{O}(N^{3})$ for both backflow types. The maximal representation ability is achieved by considering backflow terms from all sites: $Q=M$, meanwhile the complexity is $\mathcal{O}(MN^{3})$.
When generating a coefficient for the wave-function after a Lanczos optimization, the complexity is $\mathcal{O}(N^{4}Q)$ for generating a coefficient $w(\mathbf{S})$ based on Eq.(6). During the Lanczos optimization, because of the term $\langle\Psi_{p=0}|\hat{H}^{2}|\Psi_{p=0}\rangle$ when calculating the variance $\sigma_{p}^{2}$, the complexity is $\mathcal{O}(N^{5}Q)$.
III Numerical results
In this section, we present numerical results for the two-dimensional Fermi-Hubbard model achieved by the Tensor-Backflow method. To demonstrate the method’s high accuracy, we firstly present energy comparisons respect to state-of-the-art results.
The Hamiltinian of the Fermi-Hubbard model is:
$\begin{split}\hat{H}=&-t\sum_{\langle ij\rangle,\sigma}(\hat{c}_{i\sigma}^{\dagger}\hat{c}_{j\sigma}+\mathrm{H.c.})-t^{\prime}\sum_{\langle\langle ij\rangle\rangle,\sigma}(\hat{c}_{i\sigma}^{\dagger}\hat{c}_{j\sigma}+\mathrm{H.c.})\\ &+U\sum_{i}\hat{n}_{i\uparrow}\hat{n}_{i\downarrow},\end{split}$ (7)
where $t$ and $t^{\prime}$ are strengths for NN and NNN hoppings, respectively. $U$ is the strength of on-site interactions, $\hat{c}_{i\sigma}^{\dagger}(\hat{c}_{i\sigma})$ creates (destroys) a particle of spin $\sigma$ on the $i$-th site, and the particle number operator $\hat{n}_{i\sigma}=\hat{c}_{i\sigma}^{\dagger}\hat{c}_{i\sigma}$. We set tt=1 through our investigations. Energy results achieved by wave-functions before and after a Lanczos optimization are evaluated with the MC sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$.
III.1 Optimization details
In VMC optimizations, sample number and step size are important for achieving fully convergences. To identify these parameters, we have tested two optimization strategies:
  • (1)
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$ and then reducing it to $2\times 10^{-4}$;
  • (2)
    Under the sample number in the magnitude of $\mathcal{O}(4.5\times 10^{4})$, firstly setting the step size as $2\times 10^{-3}$, then increasing the sample number to $\mathcal{O}(4.5\times 10^{5})$ and maintaining the step size, finally reducing the step size to $2\times 10^{-4}$.
Figure 1: Comparisons of two optimization strategies demonstrated by nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattice under OBC, with NN backflow in wave-functions. Both strategies start from the same initial state. For strategy 1, sample number is fixed as $\mathcal{O}(4.5\times 10^{4})$, the step size is reduced from $2\times 10^{-3}$ to $2\times 10^{-4}$ at the 6000-th step. For strategy 2, sample number is increased from $\mathcal{O}(4.5\times 10^{4})$ to $\mathcal{O}(4.5\times 10^{5})$ at around 3000-th step, meanwhile maintaining the step size to $2\times 10^{-3}$. After about another 2500 steps, the step size is reduced to $2\times 10^{-4}$. Strategy 1 and 2 converge to -0.6773 and -0.6774, respectively.
Comparisons of two optimization strategies are depicted in Fig.(1). The case in the figure is nn=0.875, UU=8, $t^{\prime}$=0 on the $4\times 16$ lattice under open boundary condition (OBC), with NN backflow terms in the wave-function. From the figure, both strategies converge to similar energies. However, strategy 2 requires less optimization steps. Both strategies achieve similar ground state patterns after fully convergences, and comparisons of ground state patterns are detailed in Appendix.(A). Furthermore, we find that strategy 2 is beneficial to the convergence of all-site backflow, due to the much larger MC sample number. Therefore, strategy 2 is used for optimizing all-site backflow wave-functions. On lattices larger than 64 sites, for both optimization strategies, no more than $2\times 10^{4}$ optimization steps are used for fully convergence.
Figure 2: Energy extrapolations and comparisons respect to fPEPS Hubbard_fPEPS under OBC for the lattice of $8\times 8$(a) and $4\times 16$(b). All cases are for nn=0.875, UU=8 and $t^{\prime}$=0. Energy extrapolation for the lattice of $4\times 16$ under PBC(c) yields to the neural Pfaffian result Hubbard_NN4 . $\sigma^{2}$ is the energy variance. Red and blue solid dots depict Tensor-Backflow with NN and all-site backflow terms. For solid dots of the same color, low and high energies depict results with and without one Lanczos optimization.
III.2 Comparisons with previous results
Fig.(2)(a)(b) depict energy comparisons respect to the fPEPS method under OBC on 64-site lattices. The fPEPS is a state-of-the-art method for lattice fermions and its state representation ability can be improved by increasing the bond-dimension ($D$). From literature Hubbard_fPEPS , the complexity of fPEPS scales rapidly respect to $D$: $\mathcal{O}(D^{6})$, and the maximal bond dimension is DD=20. The fPEPS is firstly optimized by the imaginary-time-evolution method named as the simple-update (SU), followed by the gradient-optimization (GO). The fPEPS with DD=20 can achieve very high energy precision under OBC.
$n$ $U$
Lattice
Size
Boundary
Condition
$E_{p=0}$ $E_{p=1}$ Ref.
0.875 8 $8\times 8$ OBC -0.6955 -0.6982 -0.6993 Hubbard_fPEPS
$4\times 16$ -0.6797 -0.6821 -0.6830 Hubbard_fPEPS
$16\times 16$ -0.7219 -0.7257 -0.7260 Hubbard_fPEPS
$16\times 16$ PBC -0.7509 -0.7552 -0.7544 Manybody_benchmark
10 $16\times 16$ A–PBC -0.6786 -0.6828 -0.6813 Hubbard_TNS
1 8 $12\times 12$ PBC -0.5209 -0.5224 -0.5246 AFQMC_n1
A–PBC -0.5205 -0.5224 -0.5249 AFQMC_n1
44/64 4 $8\times 8$ PBC -1.1833 -1.1846 -1.1867 AFQMC_sym
80/100 $0\times 10 -1.1$086 -1.1109 -1.1147 AFQMC_sym
116/144 $2\times 12 -1.1$038 -1.1061 -1.1102 AFQMC_sym
Table 1: Energy comparisons respect to state-of-the-art results for various filling number $n$ and interaction strength $U$. All cases are for $t^{\prime}$=0. A-PBC denotes the boundary condition of antiperiodic along the x-boundary and periodic along the y-boundary. $E_{p=0}$ and $E_{p=1}$ are energies achieved by Tensor-Backflow without and with one Lanczos optimization, respectively.
In our optimizations, the initial state of Tensor-Backflow is unrestricted Hartree-Fock (UHF) state represented by a tensor of the dimension $[M,N,2]$, where $M$ is the site number $N$ is the particle number and 2 is because of spin. From the initial state, we expand the tensor’s dimension to Eq.(2) to include backflow terms tensorbackflow .
In Fig.(2)(a)(b), red and blue solid dots denote NN and all-site backflow terms, respectively. High and low energy values for same-colored solid dots denote without and with one Lanczos optimization, respectively. Therefore, a Lanczos optimization significantly improves energy precision for either NN backflow and all-site backflow. Comparing obtained energies after one Lanczos optimization and the fPEPS with DD=20, the NN backflow can achieve the energy precision competitive to the SU optimized fPEPS, meanwhile the all-site backflow can achieve the energy precision competitive to GO optimized fPEPS. The Tensor-Backflow has no limitations on boundary conditions, and energies on the $4\times 16$ lattice under PBC are depicted in Fig.(2)(c).
To obtain the energy extrapolation under the limit of zero energy variance, we linearly fit the four solid dots for each lattice size Hubbard_TNS ; lanczos . In Fig.(2), energy extrapolations are denoted by blue solid lines. From Fig.(2)(a)(b), extrapolated energies under OBC for $8\times 8$ is -0.6993, and for $4\times 16$ is -0.6826. Extrapolated energies are close to state-of-the-art energies: -0.6993 for $8\times 8$ and -0.6830 for $4\times 16$, achieved by the fPEPS with DD=20 within GO optimized. Based on results of OBC, the extrapolated energy of PBC in Fig.(2)(c) is -0.7642, matching the state-of-the-art neural Pfaffian result Hubbard_NN4 .
Energy comparisons respect to fPEPS Hubbard_fPEPS , NQS Manybody_benchmark , tensor networks Hubbard_TNS and AFQMC AFQMC_n1 ; AFQMC_sym are detailed in Tab.(1). In the table, $E_{p=0}$ is from the Tensor-Backflow and $E_{p=1}$ is from a Lanczos optimization of the converged wave-function. For OBC, reference energies are from GO optimized fPEPS with DD=20 Hubbard_fPEPS . On lattices of $8\times 8$ and $4\times 16$ under OBC, the wave-function’s type is all-site backflow. On the $16\times 16$ lattice under OBC, the wave-function’s type is NN backflow.
In this work, the relative error of energy $E$ is defined as: $(E-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. Under OBC, relative errors of $E_{p=1}$ respect to fPEPS results are $1.6\times 10^{-3}$, $1.3\times 10^{-3}$ and $4.1\times 10^{-4}$ for $8\times 8$, $4\times 16$ and $16\times 16$ lattices, respectively. Therefore, Tensor-Backflow has high precision for OBC. For boundary conditions other than OBC, from the VMC method in literature Manybody_benchmark , there is $E_{\mathrm{ref}}=-0.7544$ for nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, and $E_{p=1}$ is $1\times 10^{-3}$ lower. For nn=0.875 and UU=10, with antiperiodic along the x-boundary and periodic along the y-boundary, denoted by A-PBC in Tab.(1), the reference energy is from the energy extrapolation result of the tensor network method in literature Hubbard_TNS , meanwhile $E_{p=1}$ is $2.2\times 10^{-3}$ lower. In Tab.(1), for each reference energy from literatures AFQMC_n1 ; AFQMC_sym , $E_{p=1}$ is competitive and the relative error is in the magnitude of $\mathcal{O}(10^{-3})$.
$n/U$
Boundary
Condition
$t^{\prime}$
Lattice
Size
$E_{p=1}$ Ref. Hubbard_NN3
Relative
Error
0.875/8 PBC 0 $8\times 16$ -0.7568 -0.7568
$16\times 16$ -0.7552 -0.7563 $1.5\times 10^{-3}$
-0.2 $8\times 16$ -0.7410 -0.7419 $1.2\times 10^{-3}$
$12\times 12$ -0.7420 -0.7414 -$8.1\times 10^{-4}$
OBC 0 $16\times 16$ -0.7257 -0.7275 $2.5\times 10^{-3}$
Table 2: Energy comparisons respect to state-of-the-art NQS results Hubbard_NN3 for nn=0.875, UU=8 under various $t^{\prime}$, lattice sizes and boundary conditions, with NN backflow for $t^{\prime}$=0 and NNN backflow for $t^{\prime}$=-0.2. The relative error is defined as: $(E_{p=1}-E_{\mathrm{ref}})/|E_{\mathrm{ref}}|$. $E_{p=1}$ are competitive or even lower than state-of-the-art NQS results. In particular, The result for $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC is lower than the NQS result.
From comparisons in Tab.(1), the Tensor-Backflow can achieve competitive results respect to mainstream methods. Furthremore, we compare our results to recent state-of-the-art NQS method Hubbard_NN3 in Tab.(2). In the table, except the result on OBC (detailed in Sec.III.3), each result was achieved by the first-order-gradient optimization with the step number less than $2\times 10^{4}$, without any pinning-field appied during optimizations. However, for NQS, each optimization was performed by an improved stochastic reconfiguration method with the training procedure consisting of pre-training, applying pinning-fields and the optimization step number in the magnitude of $\mathcal{O}(10^{5})$ Hubbard_NN3 . In particular, for the case of $t^{\prime}$=-0.2, nn=0.875, UU=8 on the $12\times 12$ lattice under PBC, the achieved energy is $8.1\times 10^{-4}$ lower than the NQS result, and the optimization is performed by first-order-gradient method in less than $1\times 10^{4}$ steps, optimization procedure is detailed in Sec.III.3. Therefore, the Tensor-Backflow is efficient for optimizations meanwhile has a strong state representation ability.
Tab.(3) denotes parameter numbers used in Tensor-Backflow for various lattice sizes and backflow types, as well as energy precision respect to two state-of-the-art methods under OBC. According to Eq.(2), for the filling of nn=0.875 on 64-site lattices, the parameter number for NN backflow and all-site backflow are 286720 and 3670016, respectively. Meanwhile, the parameter number for NN backflow on the $16\times 16$ lattice is 4587520. In the table, we compare $E_{p=1}$ to results achieved by state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 . From energy comparisons in the table, NN backflows can achieve low energy relative errors in the magnitude of $\mathcal{O}(10^{-3})$. Meanwhile, considering all backflow terms can further improve energy precision.
The parameter number of Tensor-Backflow is much lower comparing to the state-of-the-art fPEPS. On the $16\times 16$ lattice under OBC, the fPEPS with DD=14 has $5.7\times 10^{6}$ parameters, and the achieved energy is -0.7224. The fPEPS with DD=16 has $9.4\times 10^{6}$ parameters, and the achieved energy is -0.7251. The fPEPS with DD=18 has $1.5\times 10^{7}$ parameters, and the achieved energy is -0.7260. The fPEPS with the maximal bond dimension DD=20 has $2.2\times 10^{7}$ parameters, with the similar energy as DD=18 Hubbard_fPEPS . For the Tensor-Backflow on the $16\times 16$ lattice, there are $E_{p=0}=$-0.7219 and $E_{p=1}=$-0.7257. Under the similar energy precision, without a Lanczos optimization, the Tensor-Backflow has many lower parameters comparing to the fPEPS with DD=14. After a Lanczos optimization, same energy precision is achieved by the Tensor-Backflow with significantly lower parameters comparing to the fPEPS with DD=18,20. Therefore, the Tensor-Backflow is an efficient state representation.
Backflow
Type
Lattice
Size
Parameter
Number
Relative
Error Hubbard_fPEPS
Relative
Error Hubbard_NN3
NN $8\times 8$ 286720 $2.3\times 10^{-3}$ $2.6\times 10^{-3}$
$4\times 16$ 286720 $3.6\times 10^{-3}$ $4.1\times 10^{-3}$
all-site $8\times 8$ 3670016 $1.6\times 10^{-3}$ $1.9\times 10^{-3}$
$4\times 16$ 3670016 $1.3\times 10^{-3}$ $1.8\times 10^{-3}$
NN $16\times 16$ 4587520 $4.5\times 10^{-4}$ $2.5\times 10^{-3}$
Table 3: The parameter number used in Tensor-Backflow for various site number and backflow types, and comparisons of energy precision respect to state-of-the-art fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 , with nn=0.875 and UU=8 under OBC. Relative errors are for $E_{p=1}$. Parameter number is based on Eq.(2), where $Q=5,M$ for NN and all-site backflows, respectively.
III.3 n=0.875
The filling of nn=0.875 is important for understanding the mechanism of superconductivity, especially when the value of $t^{\prime}$ is non-zero Hubbard_SC1 . The ground energies are highly competitive between various methods Hubbard_stripe . In this subsection, we present more details such as ground state properties and structure factors when nn=0.875 for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. We firstly investigate results achieved by Tensor-Backflow when $t^{\prime}$=0.
Figure 3: Patterns of ground states for nn=0.875, UU=8 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC(a)(b) and OBC(c)(d). All results are from pp=1 wave-functions with NN backflow. Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$.
In this work, most results from Tensor-Backflow are achieved by direct optimizations. For a direct optimization, an UHF is optimized beforehand, followed by constructing the Tensor-Backflow through expanding the tensor’s dimension to add backflow terms, meanwhile the Tensor-Backflow is optimized using either strategy from Sec.III.1. There are no pinning-fields applied during optimizations. When optimizing the UHF, the interaction strength is denoted as $U_{\mathrm{UHF}}$. Unless specifically emphasized, there is $U_{\mathrm{UHF}}=U$, and the UHF is not sufficiently optimized when $U_{\mathrm{UHF}}=U$ to avoid introducing any prior bias UHF . Futhermore, Tensor-Backflow can override properties from the initial UHF, namely, the converged wave-function is essentially a different state from the initial UHF, detailed comparisons are depicted in Appendix.(B).
Fig.(3)(a)(b) depict the linear stripe order of the ground state for UU=8 on the $16\times 16$ lattice under PBC by a direct optimization. Fig.(3)(a) depicts the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$ and Fig.(3)(b) depicts the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. As depicted by the ground state’s pattern, periods for the SDW and CDW are 16 and 8, respectively. For PBC, the ground state should be translational invariant thus an obvious pattern is not consistent with the true ground state. Since translational invariance is not enforced in the variational wave-function, even the obtained energy is competitive, the optimization is converged to a subplace of the true ground state.
Lattice
Size
Boundary
Condition
$U$
Direct
Optimization
Transfer
from UU=8
$16\times 16$ PBC 8 -0.7552
10 -0.6815 -0.6831
12 -0.6303 -0.6317
A-PBC 10 -0.6813 -0.6828
Table 4: Energy comparisons of $E_{p=1}$ between different initializations for nn=0.875 and $t^{\prime}$=0, with NN backflow in wave-functions. For the direct optimization of each case, the initial UHF is insufficiently optimized with $U_{\mathrm{UHF}}=U$. The direct optimization of UU=8 achieves the ground state of linear stripe order, and such order is preserved when it is transferred for other $U$ and boundary conditions. Energy convergences of transferring from UU=8 are detailed in Appendix.C. Energies with linear stripe orders are much lower than those by direct optimizations.
Figure 4: Charge and spin structure factors for nn=0.875 and $t^{\prime}$=0 under various $U$ and boundary conditions. All results are from pp=1 wave-functions, with NN backflow in wave-functions. (a)(b)On the $16\times 16$ lattice, results of UU=8 under PBC are obtained by a direct optimization, meanwhile other results are obtained by the initialization with the UU=8 PBC wave-function. (c)(d)Results of UU=8 under PBC by a direct optimization on the $8\times 32$ lattice.
For nn=0.875, UU=8, $t^{\prime}$=0 on the $16\times 16$ lattice under OBC, we transfer the wave-function converged under PBC as the initial state and then performed optimizations. The achieved state-of-the-art energy is -0.7257 as reported in Tab.(1) and (3). Meanwhile, Fig.(3)(c)(d) depict the average spin and charge density per site, respectively. Even the stripe order in the initial state has the same period as the ground state under OBC, the optimization can slightly modify the stripe’s pattern according to the boundary condition. Patterns of the ground state under OBC yield to fPEPS results Hubbard_fPEPS .
For other interaction strengths such as UU=10 and 12, Tab.(4) denote energy comparisons between the linear stripe order and direct optimizations, where linear stripe order is achieved when transferring the wave-function converged for UU=8 and PBC. In particular, direct optimization achieves diagonal stripe order when UU=12 under PBC. Ground state patterns of direct optimizations are depicted in Appendix.B. Energy convergences of transferring from UU=8 is detailed in Appendix.C. Energy comparisons are depicted in Tab.(4), and all energies are evaluated by pp=1 wave-functions. Because the energy of linear stripe order is much lower than other orders achieved by direct optimizations, the ground state is linear stripe order for large $U$.
Lattice
Size
$t^{\prime}$
Backflow
Type
$E_{p=0}$ $E_{p=1}$
$8\times 16$ 0 NN -0.7529 -0.7568
$8\times 24$ -0.7500 -0.7546
$8\times 32$ -0.7505 -0.7547
$8\times 16$ -0.2 -0.7354 -0.7406
$8\times 16$ NNN -0.7359 -0.7410
$12\times 12$ -0.7335 -0.7388
$16\times 16$ -0.7335 -0.7386
Table 5: Energy results obtained by direct optimizations for nn=0.875 and UU=8. For all cases in the table, initial UHFs are insuffieintly optimized with $U_{\mathrm{UHF}}=U$. For $t^{\prime}$=-0.2, NNN backflow terms are considered in the Tensor-Backflow. On the $8\times 16$ lattice, NNN backflow achieves the energy slightly lower than that comparing to NN backflow.
$n,U,t^{\prime}$
Backflow
Type
$U_{\mathrm{UHF}}$ $E_{p=0}$ $E_{p=1}$
0.875, 8, -0.2 NNN 0 -0.7344 -0.7395
4 -0.7370 -0.7420
8 -0.7335 -0.7388
Table 6: Comparisons of converged energies on the $12\times 12$ lattice from different $_{\mathrm{UHF}}, UU$$_{\mathrm{UHF}}$ is the interaction strength when optimizing the UHF. The UHF is fully converged when $U_{\mathrm{UHF}}\neq U$, otherwise the UHF is insufficiently optimized when $U_{\mathrm{UHF}}=U$. Initialization with $U_{\mathrm{UHF}}=4$ gives the lowest energy result for Tensor-Backflow.
Structure factors determine periods of the stripe order in the ground state. The charge and spin structure factors are defined as: $C(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}n_{i}n_{j}$, $S(\textbf{k})=\frac{1}{M^{2}}\sum_{ij}e^{i\textbf{k}\cdot(\textbf{r}_{i}-\textbf{r}_{j})}s_{i}^{z}s_{j}^{z}$, where $n_{i}$ and $s_{i}^{z}$ are defined in Fig.(3). $C(\textbf{k})$ is set to zero when $k_{x}=k_{y}$=0. All results of structure factors are obtained by pp=1 wave-functions. Fig.(4) depicts structure factors for the filling of nn=0.875 under various $U$ and boundary conditions, on lattices of $16\times 16$ and $8\times 32$. As depicted by Fig.(4)(a)(b), charge structure factor peaks at $\pi/4$ and spin structure factor peaks at $7\pi/8$, corresponding to periods of 8 and 16 for CDW and SDW, respectively. Fig.(4)(c)(d) depicts results of a direct optimization on the $8\times 32$ lattice for nn=0.875 and UU=8 under PBC. A vertical stripe order is obtained, meanwhile structure factors peaks at same locations as on the $16\times 16$ lattice. On the $16\times 16$ lattice, the lattice’s side length of 16 corresponds to a whole period of the SDW, thus peaks on the $16\times 16$ lattice are higher than those on the $8\times 32$ lattice.
Tab.(5) depicts additional energy results for nn=0.875 and UU=8 under PBC for both $t^{\prime}$=0 and $t^{\prime}$=-0.2. All results in the table are obtained by direct optimizations. For $t^{\prime}$=0 on the $8\times 16$ lattice, the vertical stripe order is obtained with same periods as those on the $16\times 16$ lattice.
In particular, energies for $t^{\prime}$=-0.2 are depicted in Tab.(5). The $t^{\prime}$=-0.2 is important for understanding mechanisms of superconductivity Hubbard_SC1 . From energy comparisons in Tab.(5), we find that NNN backflow (QQ=9) achieves slightly lower energy than NN backlfow (QQ=5) for the case of $t^{\prime}$=-0.2 on the $8\times 16$ lattice. Because of the non-zero NNN hopping strengths, NNN backflow are considered for all cases of $t^{\prime}$=-0.2.
Figure 5: Patterns of ground states for nn=0.875, UU=8 and $t^{\prime}$=-0.2 under PBC on the $8\times 16$ lattice, from the pp=1 wave-function with NNN backflow. Left-sided (right-sided) figure denotes the average spin (density) per site, within definitions from Fig.(3).
Figure 6: Comparisons of spin structure factors between $t^{\prime}$=0 and -0.2 for nn=0.875 and UU=8 under PBC. All results are from pp=1 wave-functions, with NN backflow for $t^{\prime}$=0 and NNN backflow for $t^{\prime}$=-0.2. Number of k-points are increased from $17^{2}$ to $65^{2}$ when calculating spin structure factors. The horizontal axis is $k_{x}$ for $t^{\prime}$=-0.2 and $k_{y}$ for $t^{\prime}$=0, meanwhile vertical axis is $S(k_{x},k_{y}=\pi)$ for $t^{\prime}$=-0.2 and $S(k_{x}=\pi,k_{y})$ for $t^{\prime}$=0.
Because cases of $t^{\prime}$=-0.2 are more challenging than those of $t^{\prime}$=0, optimizations are easily to get trapped on local minima, especially on large lattices. We show that choosing a proper initial UHF is very helpful to avoid the local minima for the final converged energy Hubbard_NN4 . In Tab.6, we compare energies converged by Tensor-Backflow with initial UHFs from different strengths of $U$ on the $12\times 12$ lattice. For cases of UU=8, $U_{\mathrm{UHF}}$=4 leads to the best converged energy of Tensor-Backflow. Ground state patterns on the $12\times 12$ lattice for various $U_{\mathrm{UHF}}$ are depicted in Appendix.D.
On the $8\times 16$ lattice, the ground state pattern for $t^{\prime}$=-0.2 is depicted in Fig.(5), with definitions of average spin and density per site from Fig.(3). For the case on the $8\times 16$ lattice, the initial UHF is insufficiently optimized under $U_{\mathrm{UHF}}=U$. From Fig.(5), a width-4 stripe is achieved, yielding to the result on the same lattice from NQS Hubbard_NN3 . In the optimization of the NQS, energies were compared between vertical and horizontal stripes and the horizontal stripe achieved lower energy Hubbard_NN3 . In this work, the horizontal stripe is achieved naturally after optimizations without pinning-fields.
In Fig.(6), we compare spin structure factors between $t^{\prime}$=0 and -0.2. To reval detailed differences, we increase the number of k-points from $17^{2}$ to $65^{2}$. Thus, there are totally 33 k-points in the x-axis. From the figure, for $t^{\prime}$=0 peaks are sharply distributed and centered on $7\pi/8$. However, the peak for $t^{\prime}$=-0.2 is wider comparing to that for $t^{\prime}$=0. The wider peak reveals competitions between stripe orders of different periods in the ground state for $t^{\prime}$=-0.2, thus the optimization is easily to get stuck on a local minima. The high optimization difficulty of $t^{\prime}$=-0.2 is depicted by Tab.(6) on the $12\times 12$ lattice. Therefore, comparisons of various $U_{\mathrm{UHF}}$ are necessary for $t^{\prime}$=-0.2 on the $16\times 16$ lattice in near future.
III.4 n=0.8 and n=0.9375
$n$ $U$
Lattice
Size
Backflow
Type
$E_{p=0}$ $E_{p=1}$
0.8 4 $10\times 20$ NN -1.1012 -1.1037
8 $10\times 20$ -0.8755 -0.8793
$10\times 10$ -0.8801 -0.8848
$10\times 10$ all-site -0.8809 -0.8853
0.9375 4 $8\times 16$ NN -0.9384 -0.9409
$8\times 24$ -0.9394 -0.9421
$8\times 32$ -0.9401 -0.9426
Table 7: Energy results for the filling nn=0.8 and nn=0.9375, with $t^{\prime}$=0 under PBC. For all cases in the table, initial UHFs are insufficiently optimized with $U_{\mathrm{UHF}}=U$. On the $10\times 10$ lattice, all-site backflow has moderate improvement comparing to NN backflow for nn=0.8 and UU=8.
Figure 7: Charge and spin structure factors on the $10\times 20$ lattice for nn=0.8 and $t^{\prime}$=0. All results are from pp=1 wave-functions with NN backflow in wave-functions. (a)(b) and on the $8\times 32$ lattice for nn=0.9375 (c)(d), under PBC. (a)(b)For nn=0.8, the spin structure factor peaks at $4\pi/5$ when UU=8, however there are no obvious peaks when UU=4. For both UU=4,8, there are no obvious peaks in the charge structure factor. (c)(d)For nn=0.9375, the spin structure factor peaks at $15\pi/16$ when UU=4, meanwhile there are no obvious peaks in the charge structure factor.
For other fillings such nn=0.8 and nn=0.9375, energy results are presented on Tab.(7). As depicted by cases of nn=0.8, UU=8 on the $10\times 10$ lattice, the improvement of all-site backflow is moderate comparing to NN backflow. Thus, for a Hamiltonian with NN hoppings, the major state representation is achieved by the NN backflow.
Fig.(7) depicts charge and spin structure factors for nn=0.8 on the $10\times 20$ lattice and for nn=0.9375 on the $8\times 32$ lattice. Based on the phase diagram via AFQMC AFQMC_phase_diagram , the critical interaction strength $U_{c}$ between un-ordered and ordered ground states increase with the doping level $h$. Here, for nn=0.8, the spin structure factor peaks at $4\pi/5$ for UU=8, meanwhile there are no obvious peaks for UU=4. Comparing charge structure factors between UU=4 and UU=8, there is the tendency of the obvious peak in the charge structure factor when further increasing $U$. For nn=0.9375, the spin structure factor peaks at $15\pi/16$, meanwhile there are no obvious peaks in the charge structure factor. Results of nn=0.8 and nn=0.9375 in Fig.(7) yield to the phase diagram via AFQMC AFQMC_phase_diagram ; AFQMC_n0.9375 .
IV Discussions and conclusions
In this work, we benchmarked the Tensor-Backflow method along with a Lanczos optimization for various fillings $n$, interaction strengths $U$, NNN hopping strengths $t^{\prime}$, lattice sizes and boundary conditions. Comparing to current state-of-the-art methods such as fPEPS Hubbard_fPEPS , NQS Hubbard_NN3 ; Hubbard_NN4 , achieved energies are competitive or even lower. In addition to energy results, we investigate ground state properties through patterns of $s_{i}^{z}$ and $n_{i}$ as well as charge and spin structure factors. When $t^{\prime}$=0, obtained state properties are consistent with other state-of-the-art methods such as fPEPS Hubbard_fPEPS and NQS Hubbard_NN3 for nn=0.875 as well as AFQMC for nn=0.8 AFQMC_phase_diagram and nn=0.9375 AFQMC_n0.9375 . For $t^{\prime}$=-0.2 on the $8\times 16$ lattice, the width-4 horizontal stripe is achieved, matching the result from NQS Hubbard_NN3 . On the $12\times 12$ lattice, achieved energy is lower than the state-of-the-art NQS result Hubbard_NN3 , meanwhile a width-4 stripe is obtained based on $n_{i}$.
Figure 8: Density-density correlations for nn=0.875, UU=8 and $t^{\prime}$=0 on the $8\times 32$ lattice under PBC. With NN backflow in the wave-function. Blue dots depict the pp=0 wave-function and yellow dots depict the pp=1 wave-function. For both pp=0 and pp=1, results are obtained based on 46080 MC samples, with the interval between two samples twice as the lattice size.
From our results, although a Lanczos optimization significantly improves the energy, the property of the state is determined by the pp=0 wave-function. Namely, a Lanczos optimization can’t override the property of the original wave-function, as patterns of $s_{i}^{z}$ and $n_{i}$ are similar between pp=0 and pp=1 wave-functions. For example, Fig.(8) depicts density-density correlations from pp=0 and pp=1 wave-functions for nn=0.875 and $U=8$ on the $8\times 32$ lattice under PBC. From the figure, the CDW with the period of 8 is already in the pp=0 wave-function and there are no significant differences for pp=1 results. Therefore, differences between correlations from pp=0 and pp=1 wave-functions are negligible.
Pair correlations determine the existence of superconductivitiy tJ_fPEPS ; pair_pair_correlation_1 ; pair_pair_correlation_2 . In this work, based on comparisons between pp=0 and pp=1 wave-functions, to avoid high calculation complexity, we evaluate $d$-wave pair correlations based on pp=0 wave-functions, with pair correlations defined in the literature tJ_fPEPS . For the case of nn=0.875 and UU=8 on the $16\times 16$ lattice under PBC, when $t^{\prime}$=0 and $t^{\prime}$=-0.2, pair correlations are in the magnitude of $\mathcal{O}(10^{-5})$ when the distance between two pairs is $L/2$, where $L$ is the side length of the square lattice. Because optimizations are easily to get trapped on local minima for cases of $t^{\prime}$=-0.2, more careful and intensive investigations are necessary in the future work.
Based on results in this work, the Tensor-Backflow is an efficient state representation for the lattice Fermi-Hubbard model. Furthermore, there are several ways to improve the energy precision. For example, considering more backflow terms comparing to hopping terms in the Hamiltonian, comparing energies under various initializations, comparing ground state patterns with and without applying pinning-fields, applying more advanced optimization methods such as the stochastic reconfiguration, etc. Currently, the wave-function is not symmetry enforced, thus translational invariance is not preserved during optimizations. For PBC, the energy precision can be further improved by enforcing translational invariance on the wave-function. The Tensor-Backflow has no restrictions on lattice shapes and boundary conditions, thus it is potentially an efficient and universal method for solving lattice fermion systems.
Figure 9: Comparisons of ground state patterns between two optimization strategies in Sec.III.1. The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $4\times 16$ lattie under OBC, with NN backflow in wave-functions. Left-sided figures denote the average spin per site: $s_{i}^{z}=(n_{i}^{\uparrow}-n_{i}^{\downarrow})/2$, and right-sided figures denote the average density per site: $n_{i}=n_{i}^{\uparrow}+n_{i}^{\downarrow}$. (a)(b) and (c)(d) are achieved by optimization strategies 1 and 2, respectively.
Figure 10: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF and (c)(d)the converged wave-function. The case is nn=0.875, UU=8 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
V Acknowledgement
X. Liang thanks useful discussions with S. Zhang, W.-Y. Liu and A. Chen. X. Liang thanks Eric J. Walter for resource allocations on the William & Mary Sciclone cluster. The author acknowledges William & Mary Research Computing and the Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the results reported within this paper. Part of numerical results were obtained on the Bridges-2 at Pittsburgh Supercomputing Center through allocation (phy240244).
Figure 11: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF, (c)(d)the converged wave-function. The case is nn=0.875, UU=10 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
Figure 12: Comparisons of structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$ from (a)(b)the initial UHF, (c)(d)the converged wave-function. The case is nn=0.875, UU=12 and $t^{\prime}$=0 on the $16\times 16$ lattice under PBC, with NN backflow in the wave-function.
Appendix A Comparing ground state patterns of two optimization strategies
Fig.(9) compares ground state patterns between two optimization strategies in Sec.III.1, and results are from pp=0 wave-functions. The case is nn=0.875, UU=8 on the $4\times 16$ lattice under OBC. From the figure, both optimizations strategies achieve similar ground state patterns. Meanwhile, ground state patterns in Fig.(9) yield to those from state-of-the-art NQS results Hubbard_NN3 .
Figure 13: Energy convergence for nn=0.875 and $t^{\prime}$=0 on the $16\times 16$ lattice when the wave-function is transferred from UU=8, with NN backflow in wave-functions. More steps are required when the case has larger difference than UU=8 and PBC. Converged energies are -0.6787, -0.6272 and -0.6786 in (a)(b) and (c), respectively.
Appendix B Comparisons between initial UHFs and subsequent Tensor-Backflow wave-functions
We show that Tensor-Backflow achieves an essentially different state from the initial UHF by structure factors $C(\mathbf{k})$ and $S(\mathbf{k})$. For nn=0.875 on the $16\times 16$ lattice under PBC, cases of UU=8,10,12 are depicted in Fig.(10), Fig.(11) and Fig.(12), respectively. Based on definitions in Sec.III.3, correlations in both $C(\mathbf{k})$ and $S(\mathbf{k})$ are classical, thus they are calculated through averages on MC samples. Here, each structure factor is obtained by 46080 MC samples, results of UHF and Tensor-Backflow are from pp=0 and pp=1 wave-functions, respectively.
In our calculations, each MC sample is flattened from two-dimensional to one-dimensional in the same way and each struture factor is calcualted by the same procedure from MC samples. Comparing (a)(b) to (c)(d) in Fig.(10), (11) and (12), properties of the converged wave-function is not simply inherited from the initial UHF. Furthermore, the Tensor-Backflow can override properties of the initial UHF. In each case, there is $U_{\mathrm{UHF}}=U$, and initial UHFs are not fully optimized in all three cases.
Appendix C Energy convergence of transferring from UU=8
Fig.(13) depicts energy convergences when transferring the converged wave-function for UU=8 and PBC to other cases. Because of the good initial state, required optimization step number is much less than that from a direct optimization. For cases denoted by (a), (b) and (c) in Fig.(13), energies achieved by direct optimizations are -0.6771, -0.6264 and -0.6771, respectively. Meanwhile, energies achieved by transferring from UU=8 are -0.6787, -0.6272 and -0.6786, respectively.
Appendix D Comparing ground state patterns for various $U_{\mathrm{UHF}}$ for $t^{\prime}$=-0.2
For the case of nn=0.875, UU=8, $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC, comparisons of ground state patterns of $U_{\mathrm{UHF}}$=0, 4 and 8 are depicted in Fig.(14), with converged energies from Tab.(6). All optimization are performed without pinning-fields applied.
From the figure, different periods of stripes are achieved with different $U_{\mathrm{UHF}}$ Hubbard_NN4 . Therefore, $U_{\mathrm{UHF}}$ is essential to avoid local minima when optimizing Tensor-Backflow. Based on the density $n_{i}$, the lowest energy of -0.7420 is achieved by the width-4 stripe, yielding to the state-of-the-art NQS result Hubbard_NN3 .
Figure 14: Ground state patterns for nn=0.875, UU=8, $t^{\prime}$=-0.2 on the $12\times 12$ lattice under PBC for various $U_{\mathrm{UHF}}$, with NNN backflow in wave-functions. Left-sided figures denote the averate spin per site: $s_{i}^{z}$ and right-sided figures denote the averate density per per site: $n_{i}$. Results of $U_{\mathrm{UHF}}$=0, 4, 8 are depicted in subfigures (a)(b), (c)(d) and (e)(f), respectively. All results are from pp=1 wave-functions. No pinning-fields are applied during optimizations.